4. Find the effective bimonthly interest rate equivalent to:

(a) nominal annual interest of 9%, compounded 6 times per year;

(b) nominal annual discount of 6%, compounded quarterly;

(c) 1/2 nominal annual interest of 8%, compounded continuously.
Answer #1

4. (a) Effective Bi-monthly rate = 0.09/6 0.015 = 1.5% (Divide by 6 since there are six bi-monthly periods)

(b) Quarterly rate = 6%. Effective annual rate (1+0.06/4)^4 -1 = 6.136% and effectvie bi-monthly rate = 6.136/6 = 1.02%

Effective bi-monthly rate = 1.02%

c) Effective annual rate = 8% = 1*e^(0.08*1) = 8.32%. So bi monthly rate = 1.386% = 1.39%

Effective bi monthly rate = 1.39%